# TwinDiffusion: Enhancing Coherence and Efficiency in Panoramic Image Generation with Diffusion Models

### v3

@@ Abstract
Diffusion models have emerged as effective tools for generating diverse and high-quality content. However, their capability in high-resolution image generation, particularly for panoramic images, still faces challenges such as visible seams and incoherent transitions. In this paper, we propose TwinDiffusion, an optimized framework designed to address these challenges through two key innovations: Crop Fusion for quality enhancement and Cross Sampling for efficiency optimization. We introduce a training-free optimizing stage to refine the similarity of the adjacent image areas, as well as an interleaving sampling strategy to yield dynamic patches during the cropping process. A comprehensive evaluation is conducted to compare TwinDiffusion with the existing methods, considering factors including coherence, fidelity, compatibility, and efficiency. The results demonstrate the superior performance of our approach in generating seamless and coherent panoramas, setting a new standard in quality and efficiency for panoramic image generation.
@@ 1 Introduction
Over the past few years, diffusion models [39, 16, 27] have demonstrated their creativity in generation tasks. They define a pair of forward and reverse Markov chains to learn the data distribution, which bypasses the limitations in other types of generative models like GANs [7, 4], VAEs [23, 6] and Flows [19]. Diffusion models have risen as effective tools for broad applications, spanning from high-quality images to multi-type content creation [14]. With the growing interest in full information records, immersive virtual reality, artistic expression, and historical preservation, the synthesis of long scrolls is gaining more and more attention, particularly in panoramic image generation tasks [2, 11, 40].
Recent advancements have exhibited the expansibility of pre-trained diffusion models in generating images of arbitrary dimensions, such as super-resolution diffusion [13, 28, 27] and area fusion strategies [38, 41, 18]. The latter often involves cropping from a large space into small patches for individual processing, as well as conducting specific guidance to fuse them together, providing more controllability than the former. However, achieving crop-wise high-resolution generation is non-trivial. To our knowledge, MultiDiffusion [2] represents the state-of-the-art framework among the existing methods, yet it still fails to capture the relationships between neighboring image areas, resulting in unnatural connections or even visible seams in panoramas. Although a finer cropping stride could ease such problems, it comes with a higher time cost.
@@
  • Performance Trade-off: Moreover, we analyze the key factors in TwinDiffusion that impact its performance, including the timestep for introducing the crop optimizing stage, the Lagrange multiplier in Crop Fusion loss function, the view stride and the cross stride for sampling image patches. This thorough discussion gives insights into the condition of quality-efficiency balance with our method.
@@ 2 Related Work
Diffusion models [32] are inspired by non-equilibrium thermodynamics. They define two Markov chains for forward and reverse processes, namely diffusion and denoising. The forward process is to perturb a data distribution $x_{0}\sim q(x)$ into a standard Gaussian distribution $x_{T}\sim\mathcal{N}(0,I)$ with $T$ steps of noise injection. This process is reversed to recreate the sample $x_{0}$ that obeys the original data distribution from a Gaussian noise input. Specifically, the reverse process involves training a network to approximate $q(x_{t-1}\mid x_{t})$, and then sampling from $\mathcal{N}(0,I)$ iteratively with the trained model. From DDPM [16] to DDIM [33] to LDM [27], diffusion models have paved the way for text-to-image generation, boosting AI-painting applications like Stable Diffusion [27] and DALLE2 [26]. By capturing the spatial-temporal distribution features, they also show promising prospects in generating multi-modal contents such as videos [3, 10, 21], audio [17, 20], and 3D objects [5, 36, 34].
@@
Extensive studies have been dedicated to leveraging diffusion models for controllable high-resolution image generation tasks. The existing methodologies can be divided into two branches: (i) Methods that focus on super-resolution techniques utilizing diffusion processes to infer the missing information [13, 37, 28]. These works often require training on the specific datasets, combining initial noise with low-resolution images as the input to the network, which lacks portability while also imposing computational demands. (ii) Methods that center around crop-based fusion strategies within diffusion paths [2, 1, 18, 41]. These approaches offer greater flexibility by manipulating the generation process in different cropped spaces, and reconciling them in a training-free or fine-tuning manner. Among them, the MultiDiffusion framework proves to be feasible and resultful. However, its loss function only pays attention to the overlapping regions of image crops, ignoring the non-overlapping adjacent areas, which reduces itself into a naive weighted mean method. Hence, we see room for improvement.
@@ 3 Method
Figure 2: Illustration of the generation process when applying our approach to a panorama. The process begins with the mapping function $F_{i}$ transforming the image crops into the panoramic space $Z^{\prime}$. This results in a sequence of overlapping crops $\mathbf{x}_{t}^{1},\mathbf{x}_{t}^{2},\ldots,\mathbf{x}_{t}^{n}$ arranged spatially, each having an independent denoising path. Our goal is to optimize $\mathbf{X}_{t}^{i}$ within the constraints of its adjacent neighbor and itself as well, thus ensuring a unified and progressive fusion of crops. To achieve this alignment, our objective function Eq. 4 is defined into two mutual-restricted parts and reaches the minimizer ${\mathbf{x}_{t}^{i}}^{*}$ in each denoising timestep: (i) The matching term: differences at the overlaps of $\mathbf{x}_{t}^{i}$ and its neighbor $\mathbf{x}_{t}^{i-1}$. (ii) The regularization term: deviations between ${\mathbf{x}_{t}^{i}}^{*}$ and its unoptimized self $f_{\Phi}({\mathbf{x}_{t+1}^{i}}^{*},t,C)$.
@@ 3.1 Preliminary
To start with, we introduce a pre-trained diffusion model denoted by $\Phi$, operating in a latent space $Z=\mathbb{R}^{h\times w\times c}$ and a textual condition $C$. Employing the deterministic DDIM sampling:
@@
Our intention is to extend $\Phi$ as a reference model to generate images in a larger space $Z^{\prime}=\mathbb{R}^{h^{\prime}\times w^{\prime}\times c^{\prime}}$, where $h^{\prime}>h$ and $w^{\prime}>w$. This can be achieved with the MultiDiffusion framework, represented by a function called the MultiDiffuser $\Psi$. It defines a set of mappings between two model spaces by:
@@
Specifically, $F_{i}$ refers to cropping the $i$-th image patch from space $Z^{\prime}$.
During the MultiDiffusion process, firstly, each crop is simultaneously and independently denoised with $f_{\Phi}:z_{t-1}^{i}=f_{\Phi}(z_{t}^{i},t,C)$ suggested by Eq. 1. Then, based on the Manifold Hypothesis, a least-square optimization for global fusion is formulated to minimize the discrepancy between each crop $F_{i}(z^{\prime}_{t})$ and its denoised counterpart $f_{\Phi}(F_{i}(z^{\prime}_{t}),t,C)$, merging different crops into one large image $z^{\prime}$. According to the properties of $\Psi$, its optimization problem has an analytical solution. Thus, the minimizer $z^{\prime}$ turns out to be a weighted average value:
$z^{\prime}_{t}=\frac{\sum_{i}F_{i}(z^{\prime}_{t})}{\sum_{i}W_{i}}$ (3)
with $W_{i}$ represents the weight matrix of the $i$-th crop.
@@ 3.2 Image Crop Fusion
At each denoising timestep $t$, we get a spatial-ordered sequence of overlapping crops $\mathbf{x}_{t}^{1},\mathbf{x}_{t}^{2},\ldots,\mathbf{x}_{t}^{n}$ generated by the mapping function Eq. 2 in a panoramic space $Z^{\prime}$. For each crop $\mathbf{x}_{t}^{i}$, our goal is to align its overlapping part of the adjacent crop $\mathbf{x}_{t}^{i-1}$ as closely as possible, while limiting the deviation from the crop itself. Thus we present:
${\mathbf{x}_{t}^{i}}^{*}$ $=\arg\min_{\mathbf{x}_{t}^{i}\in Z^{\prime}}\mathcal{L}(\mathbf{x}_{t}^{i}\mid\mathbf{x}_{t}^{i-1},{\mathbf{x}_{t+1}^{i}}^{*},C)$ (4)
$=\arg\min_{\mathbf{x}_{t}^{i}\in Z^{\prime}}\parallel M_{r}^{i-1}\odot\mathbf{x}_{t}^{i-1}-M_{l}^{i}\odot\mathbf{x}_{t}^{i}\parallel^{2}+$
$\quad\quad\quad\quad\quad\;\parallel f_{\Phi}({\mathbf{x}_{t+1}^{i}}^{*},t,C)-\mathbf{x}_{t}^{i}\parallel^{2}$
as our optimization task, where $\mathbf{x}^{*}$ denotes the optimized crop, $M_{l},M_{r}\in\{0,1\}^{h\times w}$ represent the binary masks covering the crop’s left and right overlapping regions, and $\odot$ is the Hadamard product.
The second part of Eq. 4 serves as a regularization term, therefore, the objective function of TwinDiffusion can be formulated as follows:
min $\parallel M_{r}^{i-1}\odot\mathbf{x}_{t}^{i-1}-M_{l}^{i}\odot\mathbf{x}_{t}^{i}\parallel^{2}$ (5)
s.t. $\parallel f_{\Phi}({\mathbf{x}_{t+1}^{i}}^{*},t,C)-\mathbf{x}_{t}^{i}\parallel^{2}\enspace\leq\;\aleph$
@@
$L(\mathbf{x}_{t}^{i},\lambda)$ $=\parallel M_{r}^{i-1}\odot\mathbf{x}_{t}^{i-1}-M_{l}^{i}\odot\mathbf{x}_{t}^{i}\parallel^{2}+$ (6)
$\lambda(\parallel f_{\Phi}({\mathbf{x}_{t+1}^{i}}^{*},t,C)-\mathbf{x}_{t}^{i}\parallel^{2}-\;\aleph)$
where $\lambda$ is the Lagrange multiplier associated with the constraints of adjacent but non-overlapping regions in the panorama. Using the Karush-Kuhn-Tucker (KKT) conditions, we can reach an optimal solution, which demonstrates that our Crop Fusion is a training-free method with closed-form optimization.
@@ 3.3 Cross Sampling
$z_{t}^{i}=\mathcal{F}_{i}^{(k)}(z_{t}^{\prime}),\quad k=t\!\!\!\mod\>c$ (7)
where $c$ controls the interleaving frequency and $k$ denotes the sampling mode. Staggering in $c$ times, our method dynamically forms a set $\mathcal{X}=\{\mathbf{x}_{i,j}\mid 1\leq i\leq c,1\leq j\leq n\}$ consists of $c$ groups, each containing $n$ overlapping crops, with spatial locations incrementing by the cross stride $s_{c}$. Then, in $T$ rounds of denoising, we alternate between using these crop groups for sampling in $c$ staggered spaces.
By incorporating the Cross Sampling method, TwinDiffusion takes more flexible control over the fine degree of overlap between neighboring image crops, filling the gaps caused by the enlarged view stride. Our experimental results in Sec. 4.3 have demonstrated that with this straightforward yet effective solution, we can double $s_{v}$ to cut the generation time in half or more (with a larger $s_{v}$), while maintaining panoramic image quality on par with the original MultiDiffusion.
@@ 4.1 Panorama with Twin Crops
As illustrated in Fig. 3 (top), our method generates a pair of images that locally resemble each other like twins.$I_{1}$ and $I_{2}$ represent the first and second images respectively, generated from latent noise satisfying $M_{r}^{1}\odot z_{1}=M_{l}^{2}\odot z_{2}$. $I_{2}^{*}$ corresponds to $I_{2}$ with our optimization. It successfully retains the content of original $[I_{2}]_{r}$ while closely aligning its left part to $[I_{1}]_{r}$, achieving a seamless fusion of $[I_{1}]_{r}$ and $[I_{2}]_{l}$.
We also conduct a stress test on TwinDiffusion’s ability to fuse image crops. Specifically, we replace $f_{\Phi}({\mathbf{x}_{t+1}^{i}}^{*},t,C)$ in the regularization term of Eq. 5 with a constant reference $\tilde{\mathbf{x}}_{t}^{i}$, representing the raw $\mathbf{x}_{t}^{i}$ following its unoptimized denoising trajectory. The results in Fig. 3 (bottom) demonstrate that even under such conditions, our method still achieves a natural and seamless crop fusion. This work ensures higher consistency in the subsequent panorama generation.
Figure 3: Applying our Crop Fusion method to generate twin images. Top: the optimized $I_{2}^{*}$ exhibits a seamless fusion effect that meets our expectations. Bottom: We further test its limits by fixing the regularization term of Eq. 5, and the results demonstrate the robustness of our method under extreme conditions.
@@
Beyond a single pair of images, we generalize this approach to a sequence of images, i.e., crops $\mathbf{x}_{t}^{1},\mathbf{x}_{t}^{2},\ldots,\mathbf{x}_{t}^{n}$ within panoramas. Our optimization is applied to each pair of neighboring crops, promoting a high degree of similarity between $[\mathbf{x}_{t}^{i}]_{r}$ and $[\mathbf{x}_{t}^{i+1}]_{l}$, thereby resulting in higher-quality panoramas with consecutive twin crops as depicted in Fig. 1. Refer to Appendix A for more implementation examples.

### v4

@@ Abstract
Diffusion models have emerged as effective tools for generating diverse and high-quality content. However, their capability in high-resolution image generation, particularly for panoramic images, still faces challenges such as visible seams and incoherent transitions. In this paper, we propose TwinDiffusion, an optimized framework designed to address these challenges through two key innovations: the Crop Fusion for quality enhancement and the Cross Sampling for efficiency optimization. We introduce a training-free optimizing stage to refine the similarity of adjacent image areas, as well as an interleaving sampling strategy to yield dynamic patches during the cropping process. A comprehensive evaluation is conducted to compare TwinDiffusion with the prior works, considering factors including coherence, fidelity, compatibility, and efficiency. The results demonstrate the superior performance of our approach in generating seamless and coherent panoramas, setting a new standard in quality and efficiency for panoramic image generation.
@@ 1 Introduction
Over the past few years, diffusion models [15, 25, 36] have demonstrated their creativity in generation tasks. They define a pair of forward and reverse Markov chains to learn the data distribution, which bypasses the limitations in other types of generative models like GANs [4, 7], VAEs [6, 21] and Flows [18]. Diffusion models have risen as effective tools for broad applications, spanning from high-quality images to multi-type content creation. With the growing interest in full information records, immersive virtual reality, artistic expression, and historical preservation, the synthesis of long scrolls is gaining more and more attention, particularly in panoramic image generation tasks [2, 11, 38].
Recent advancements have exhibited the expansibility of pre-trained diffusion models in generating images of arbitrary dimensions, such as super-resolution diffusion [13, 34] and area fusion strategies [17, 35, 39]. The latter often involves cropping from a large space into small patches for individual processing, as well as conducting specific guidance to fuse them together, providing more controllability than the former. However, achieving crop-wise high-resolution generation is non-trivial. To our knowledge, MultiDiffusion [2] represents the state-of-the-art framework among the existing methods, yet it still fails to capture the relationships between neighboring image areas, resulting in unnatural connections or even visible seams in panoramas. Although a finer cropping stride could ease such problems, it comes with a higher time cost.
@@
  • Performance Trade-off: Moreover, we analyze the key factors in TwinDiffusion that impact its performance, including the timestep for introducing the crop optimizing stage, the Lagrange multiplier in our core function, the view stride and the cross stride for sampling image patches. This thorough discussion gives insights into the condition of quality-efficiency balance with our method.
@@ 2 Related Work
Diffusion models are inspired by non-equilibrium thermodynamics [29]. They define two Markov chains for forward and reverse processes, namely diffusion and denoising. The forward process is to perturb a data distribution $x_{0}\sim q(x)$ into a standard Gaussian distribution $x_{T}\sim\mathcal{N}(0,I)$ with $T$ steps of noise injection. This process is reversed to recreate the sample $x_{0}$ that obeys the original data distribution from a Gaussian noise input. Specifically, the reverse process involves training a network to approximate $q(x_{t-1}\mid x_{t})$, and then sampling from $\mathcal{N}(0,I)$ iteratively with the trained model. From DDPM [15] to DDIM [30] to LDM [25], diffusion models have paved the way for text-to-image generation, boosting AI-painting applications like Stable Diffusion [25] and DALLE2 [24]. By capturing the spatial-temporal distribution features, they also show promising prospects in generating multi-modal contents such as videos [3, 5, 10], audio [16, 19], and 3D objects [31, 33, 37].
@@
Extensive studies have been dedicated to leveraging diffusion models for controllable high-resolution image generation tasks. The existing methodologies can be divided into two branches: (i) methods that focus on super-resolution [13, 34] or inpainting [1] techniques utilizing diffusion processes to infer the missing information, and (ii) methods that center around crop-based fusion strategies within diffusion paths [2, 17, 39]. The former often requires training on specific datasets, combining initial noise with low-resolution images as input to the network. Additionally, they involve resizing the input images, which lacks portability and imposes computational demands. On the other hand, the latter offers greater flexibility by manipulating the generation process in different cropped spaces, and reconciling them in a training-free or fine-tuning manner. Among them, the MultiDiffusion framework proves to be feasible and resultful. However, its optimization function only pays attention to the overlapping regions of image crops, ignoring the non-overlapping adjacent areas, which reduces itself into a naive weighted mean method. Hence, we see room for improvement.
@@ 3 Method
Figure 2: Illustration of our approach applied to panorama generation. The process begins with the mapping function $F_{i}$ transforming the image crops into the panoramic space $Z^{\prime}$. This results in a sequence of overlapping crops $z_{t}^{1},z_{t}^{2},\ldots,z_{t}^{n}$ arranged spatially, each having an independent denoising path. Our goal is to optimize $z_{t}^{i}$ within the constraints of its adjacent neighbor and itself as well, thus ensuring a unified and progressive fusion of crops. To achieve this alignment, our objective function Eq. 4 is defined into two mutual-restricted parts and reaches the minimizer ${z_{t}^{i}}^{*}$ in each denoising timestep: (i) the matching term: differences at the overlaps of $z_{t}^{i}$ and its neighbor $z_{t}^{i-1}$, (ii) the regularization term: deviations between ${z_{t}^{i}}^{*}$ and its unoptimized self $f_{\Phi}(F_{i}(z^{\prime}_{t+1}),t,C)$.
@@ 3.1 Preliminary
To start with, we introduce a pre-trained diffusion model denoted by $\Phi$, operating in a latent space $Z=\mathbb{R}^{h\times w\times c}$ and a textual condition $C$. Employing the deterministic DDIM sampling [30]:
@@
Our intention is to extend $\Phi$ as a reference model to generate images in a larger space $Z^{\prime}=\mathbb{R}^{h^{\prime}\times w^{\prime}\times c}$, where $h^{\prime}>h$ and $w^{\prime}>w$. This can be achieved with the MultiDiffusion framework [2], represented by a function called the MultiDiffuser $\Psi$. It defines a set of mappings between two model spaces by:
@@
Specifically, $F_{i}$ refers to cropping the $i$-th image patch from space $Z^{\prime}$ with the stride of $s_{v}$.
During the MultiDiffusion process, firstly, each crop is simultaneously and independently denoised with $f_{\Phi}$: $z_{t-1}^{i}=f_{\Phi}(z_{t}^{i},t,C)$ suggested by Eq. 1. Then, based on the Manifold Hypothesis, a least-square optimization for global fusion is formulated to minimize the discrepancy between each crop $F_{i}(z^{\prime}_{t-1})$ and its denoised counterpart $f_{\Phi}(F_{i}(z^{\prime}_{t}),t,C)$, merging different crops into one large image $z^{\prime}$. According to the properties of $\Psi$, its optimization problem has an analytical solution. Thus, the minimizer $z^{\prime}$ turns out to be a weighted average value:
$z^{\prime}_{t-1}=\frac{\sum_{i}W_{i}\odot F_{i}^{-1}(f_{\Phi}(z_{t}^{i},t,C))}{\sum_{i}W_{i}}$ (3)
with $W_{i}$ represents the pixel weight matrix of the $i$-th crop.
@@ 3.2 Image Crop Fusion
At each denoising timestep $t$, we get a spatial-ordered sequence of overlapping crops $z_{t}^{1},z_{t}^{2},\ldots,z_{t}^{n}$ generated by the mapping function Eq. 2 in a panoramic space $Z^{\prime}$. For each crop $z_{t}^{i}$, our goal is to align its overlapping part of the adjacent crop $z_{t}^{i-1}$ as closely as possible, while limiting the deviation from the crop itself. Thus we present:
${z_{t}^{i}}^{*}$ $=\arg\min_{z_{t}^{i}\in Z^{\prime}}\mathcal{L}(z_{t}^{i}\mid z_{t}^{i-1},F_{i}(z^{\prime}_{t+1}),C)$ (4)
$=\arg\min_{z_{t}^{i}\in Z^{\prime}}\parallel M_{r}^{i-1}\odot z_{t}^{i-1}-M_{l}^{i}\odot z_{t}^{i}\parallel^{2}+$
$\quad\quad\quad\quad\quad\;\parallel f_{\Phi}(F_{i}(z^{\prime}_{t+1}),t,C)-z_{t}^{i}\parallel^{2}$
as our optimization task, where $z^{*}$ denotes the optimized crop, $M_{l},M_{r}\in\{0,1\}^{h\times w}$ represent the binary masks covering the crop’s left and right overlapping regions according to $s_{v}$, and $\odot$ is the Hadamard product. The second part of Eq. 4 serves as a regularization term, which is used to coordinate the alignment behavior of crops. Therefore, the objective function of TwinDiffusion can be formulated as follows:
min $\parallel M_{r}^{i-1}\odot z_{t}^{i-1}-M_{l}^{i}\odot z_{t}^{i}\parallel^{2}$ (5)
s.t. $\parallel f_{\Phi}(F_{i}(z^{\prime}_{t+1}),t,C)-z_{t}^{i}\parallel^{2}\enspace\leq\;\aleph$
@@
$L(z_{t}^{i},\lambda)$ $=\parallel M_{r}^{i-1}\odot z_{t}^{i-1}-M_{l}^{i}\odot z_{t}^{i}\parallel^{2}+$ (6)
$\lambda(\parallel f_{\Phi}(F_{i}(z^{\prime}_{t+1}),t,C)-z_{t}^{i}\parallel^{2}-\;\aleph)$
where $\lambda$ is the Lagrange multiplier associated with the constraints of adjacent but non-overlapping regions in the panorama. Using the Karush-Kuhn-Tucker (KKT) conditions, we can reach an optimal solution:
${z_{t}^{i}}^{*}=\begin{cases}M_{l}^{i}\odot{z_{t}^{i}}^{*}=(1+\lambda)^{-1}[M_{r}^{i-1}\odot z_{t}^{i-1}+\\ \quad\qquad\qquad\qquad\qquad\lambda M_{l}^{i}\odot f_{\Phi}(F_{i}(z^{\prime}_{t+1}),t,C)]\\[5.0pt] M_{r}^{i}\odot{z_{t}^{i}}^{*}=M_{r}^{i}\odot f_{\Phi}(F_{i}(z^{\prime}_{t+1}),t,C)\end{cases}$ (7)
which demonstrates that our Crop Fusion is a training-free method with closed-form optimization.
@@ 3.3 Cross Sampling
$z_{t}^{i}=F_{i}^{(k)}(z_{t}^{\prime}),\quad k=t\!\!\!\mod\>r$ (8)
where $r$ controls the interleaving frequency and $k$ denotes the sampling mode. Staggering in $r$ times, our method dynamically forms a set $\mathcal{Z}=\{z_{i,j}\mid 1\leq i\leq r,1\leq j\leq n\}$ consists of $r$ groups, each containing $n$ overlapping crops, with spatial locations incrementing by the cross stride $s_{r}$. Then, in $T$ rounds of denoising, we alternate between using these crop groups for sampling in $r$ staggered spaces.
Panorama seams mostly occur where the crops meet. Thus by incorporating the Cross Sampling method, TwinDiffusion takes more flexible control over the fine degree of overlap between neighboring image crops, filling the gaps caused by the enlarged view stride. Our experimental results in Sec. 4.3 have demonstrated that with this straightforward yet effective solution, we can double $s_{v}$ to cut the generation time in half or more (with a larger $s_{v}$), while maintaining panoramic image quality on par with the original MultiDiffusion.
@@ 4.1 Panorama with Twin Crops
Figure 3: Applying our Crop Fusion method to generate twin images. Top: the optimized $I_{2}^{*}$ exhibits a seamless fusion effect that meets our expectations. Bottom: We further test its limits by fixing the regularization term of Eq. 5. The results demonstrate our method’s robustness under extreme conditions.
@@
As illustrated in Fig. 3 (top), our method generates a pair of images that locally resemble each other like twins.$I_{1}$ and $I_{2}$ represent the first and second images respectively, generated from initial latent noise satisfying $M_{r}^{1}\odot z_{1}=M_{l}^{2}\odot z_{2}$. $I_{2}^{*}$ corresponds to $I_{2}$ with our optimization. It successfully retains the content of original $[I_{2}]_{r}$ while closely aligning its left part to $[I_{1}]_{r}$, achieving a seamless fusion of $[I_{1}]_{r}$ and $[I_{2}]_{l}$.
We also conduct a stress test on TwinDiffusion’s ability to fuse image crops. Specifically, we replace $f_{\Phi}(F_{i}(z^{\prime}_{t+1}),t,C)$ in the regularization term of Eq. 5 with a constant reference $\tilde{z}_{t}^{i}$, representing the raw $z_{t}^{i}$ following its unoptimized denoising trajectory. The results in Fig. 3 (bottom) demonstrate that even under such conditions, our method still achieves a natural and seamless crop fusion. This work serves as an initial validation to ensure higher consistency in the subsequent panorama generation.
@@
Beyond a single pair of images, we generalize this approach to a sequence of images, i.e., crops $z_{t}^{1},z_{t}^{2},\ldots,z_{t}^{n}$ within panoramas. Our optimization is applied to each pair of neighboring crops, promoting a high degree of similarity between $[z_{t}^{i}]_{r}$ and $[z_{t}^{i+1}]_{l}$, thereby resulting in higher-quality panoramas with consecutive twin crops as depicted in Fig. 1. Refer to Appendix A for more implementation examples.